OpenStudy (anonymous):

write the polynomial as the product of factors that are irreducible over the rationals x^4-3x^3-x^2-12x-20

OpenStudy (anonymous):

i JUST answered this question x^4 - 3x^3 - x^2 - 12x - 20 = (x^4 - x^2 - 20) - (3x^3 + 12x) = (x^2 - 5)(x^2 + 4) - 3x(x^2 + 4) = (x^2 + 4)(x^2 - 3x - 5)